OpenStudy (anonymous1):

I have a question about a derivative proof with logarithm; details in the reply. I have a question about a derivative proof with logarithm; details in the reply. @Mathematics

OpenStudy (anonymous1):

The following inequality is given: \[1 - \frac{1}{x} < \ln x < x - 1\] for every x > 0 and different from 1. Then the exercise asks to show the following: \[\lim_{x \to 0} \frac{\ln(1+x)}{x} = 1 \] but I have no idea on how to start. Can anyone give some hint? I could start from the definition of limit: \[ \left |\frac{\ln(1+x)}{x} - 1\right | < \epsilon \] whenever \[|x - 0| < \delta\] but I don't know how to continue.

OpenStudy (anonymous):

replace x by 1+x and use squeeze i think gets it in one step

OpenStudy (anonymous):

of course i could be wrong, so let me write it out and see if it works

OpenStudy (anonymous1):

\[\frac{x}{x+1}<\ln(x+1)<x\]

OpenStudy (anonymous1):

In this case, the limits of both \[\frac{x}{x+1}\] and \[x\] as x approaches zero are 0 respectively.

OpenStudy (anonymous):

yup it works. divide by x get \[\frac{1}{x+1}<\frac{\ln(x+1)}{x}<1\]

OpenStudy (anonymous1):

This is true, very interesting. Thank you.

OpenStudy (anonymous):

yw